# Language Models can Exploit Cross-Task In-context Learning for Data-Scarce Novel Tasks

### v1

@@ Abstract
**footnotetext: These authors contributed equally.
@@ 1 Introduction
In this work, we design a cross-task prompting setup (Section 2) using three different LLMs: LLaMA-2 7B and 13B, and GPT 3.5; we select $50$ different pairs of tasks where one serves as a source (i.e., context example task) and the other as target. Despite no examples from the target task presented in the context, LLMs can produce a staggering improvement over the zero-shot regime; on average, cross-task prompting improves performance by $107\%$ for LLaMA-2 7B, $18.6\%$ for LLaMA-2 13B and $3.2\%$ for GPT 3.5 (Section 1). With multiple source tasks, cross-task performance is even better than, if not comparable, usual in-task prompting (Contribution #1). However, learning from examples of different tasks is heavily sensitive to the choice of source task for a given target and the LLM is prone to copy the label space of the source task into the target. To circumvent this, we propose a pseudo-labeling based approach: in a data-scarce setup, cross-task prompting with majority voting is first employed to generate noisy, in-task examples; these are subsequently used for standard few-shot prompting (Contribution #2). Finally, we provide introductory analysis towards interpreting cross-task signal transfer by dissecting the model activations. We find that the cross-task signal transfer is abrupt and happens at later layers, with the effective layers widely varying for different target tasks (Contribution #3). In a nutshell, this is the first exploration of LLMs’ ability to learn to solve novel tasks based on the contextual signals of different task examples.
@@ 2 Prompting Techniques
Next, given two datasets $D_{s}$ and $D_{t}$ corresponding to two different tasks with task definitions $d_{s}$ and $d_{t}$, respectively, we formalize the cross-task prompting as inferring the output $\hat{y}_{t}$ for an input $x_{t}\in D_{t}$ conditioned upon a demonstration from $D_{s}$:
@@
In this setup, we denote $D_{s}$ and $D_{t}$ as source and target task datasets, respectively. Note that the formalization is equivalent to 1-shot prompting where the only provided example input-output pair comes from a different dataset. Table 13 of Appendix E presents illustrative examples of prompts in this setup. Our experiments suggest that $>$1-shot setup in cross-task prompting does not improve (and often deteriorates) performance.
@@
Examples of prompts are provided in Table 14 of Appendix E.
@@ 3 Results and Analysis
Datasets and experimental setup. Our corpus of tasks consists of ten source and five target tasks. We consider ARC-Easy Clark et al. (2018), AG-news Zhang et al. (2015), BoolQ Clark et al. (2019), Commonsense-QA Talmor et al. (2019), Conll2003-POS Tjong Kim Sang and De Meulder (2003), Conll2003-NER Tjong Kim Sang and De Meulder (2003), MNLI Williams et al. (2018), QQP Sharma et al. (2019), RACE Lai et al. (2017) and SST2 Socher et al. (2013) to be our source tasks. Our motivation remains to incorporate domain diversity and difficulty of the problems while choosing target tasks to emulate the “novel task” phenomenon as closely as possible since learning from cross-task examples makes sense only when the target task is truly data-scarce. Our target tasks are Social-i-QA Sap et al. (2019), SciQ Auer et al. (2023), MedMCQA Pal et al. (2022), Financial-Phrasebank Malo et al. (2014), and ARC-Challenge Clark et al. (2018); first four of these require domain expertise, while the last one is a more challenging version of the one present in the source tasks. In total, this gives us $50$ unique cross-task setups (see Appendix A for dataset details. Table 15 and Table 16 of Appendix E show task definitions corresponding to source and target tasks, respectively). The dataset size is standardised by sampling $10,000$ and $500$ examples for each source and target task, i.e., $|D_{s}|=10,000$ and $|D_{t}|=500$. For our experiments, we used the 7-billion and 13-billion variants of LLaMA-2 Touvron et al. (2023), and text-davinci-003 Brown et al. (2020) referred to as GPT3.5. We experiment with greedy and force decoding setup and selected greedy decoding as our standard for all experiments (see Appendix C for more details). We set the number of examples used to create cross-task context as one for all our experiments, unless mentioned explicitly.
@@
On average, cross-task prompting improves performance by 107% for LLaMA-2 7B, 18.6% for LLaMA-2 13B and 3.2% for GPT3.5. This is a strong argument for prompting in a cross-task manner when we lack labeled target data, especially using small models, which have poor zero-shot abilities  (Wei et al., 2022).
Importance of source task definitions. To further investigate the role of task definitions in cross-task prompting, we check the effect of removing source task definitions on performance. Table 12 reports the average accuracy for all possible source-target pairs corresponding to a task; We note an average drop of 11% for LLaMA-2 7B and 8% for LLaMA-2 13B when we prompt without using source task definitions (see Table 12 in Appendix). Hence, definitions play a crucial role in cross-task prompting.
@@
Semantically similar vs random example selection. As shown in Table 2, in a cross-task setup, choosing the examples randomly leads to substantially poorer performance than when we generate the context $C$ with semantically similar examples; concerning is the fact that, in many cases, it causes 0% accuracy. This may be caused by the model getting confused without any semantic alignment in the prompt. Furthermore, random labeling of in-context examples (Wei et al., 2023) result in near-random performance (see Table 11 in Appendix D).
@@ Appendix B Semantic Context creation
Appendix C Force Decoding
@@
Appendix D Random Labeling
@@
Appendix E Prompt details
@@
Appendix F Error analysis
@@
Appendix G Activation analysis on LLaMA-2 7B
Full correlation analysis is presented in Table 18 and Figure 5.

### v3

@@ Abstract
**footnotetext: Equal contribution named in alphabetical order
@@ 1 Introduction
In this work [^1] Code available at https://github.com/C-anwoy/Cross-Task-ICL, we design a cross-task prompting setup (Section 2) using three different LLMs: LLaMA-2 7B and 13B, and GPT 3.5; we select $50$ different pairs of tasks where one serves as a source (i.e., context example task) and the other as target. Despite no examples from the target task presented in the context, LLMs can produce a staggering improvement over the zero-shot regime; on average, cross-task prompting improves performance by $107\%$ for LLaMA-2 7B, $18.6\%$ for LLaMA-2 13B and $3.2\%$ for GPT 3.5 (Section 1). With multiple source tasks, cross-task performance is even better than, if not comparable, usual in-task prompting (Contribution #1). However, learning from examples of different tasks is heavily sensitive to the choice of source task for a given target and the LLM is prone to copy the label space of the source task into the target. To circumvent this, we propose a pseudo-labeling based approach: in a data-scarce setup, cross-task prompting with majority voting is first employed to generate noisy, in-task examples; these are subsequently used for standard few-shot prompting (Contribution #2). Finally, we provide introductory analysis towards interpreting cross-task signal transfer by dissecting the model activations. We find that the cross-task signal transfer is abrupt and happens at later layers, with the effective layers widely varying for different target tasks (Contribution #3). In a nutshell, this is the first exploration of LLMs’ ability to learn to solve novel tasks based on the contextual signals of different task examples.
@@ 2 Prompting techniques
Next, given two datasets $D_{s}$ and $D_{t}$ corresponding to two different tasks with task definitions $d_{s}$ and $d_{t}$, respectively, we formalize the cross-task prompting as inferring the output $\hat{y}_{t}$ for an input $x_{t}\in D_{t}$ conditioned upon a demonstration from dataset $D_{s}$:
@@
In this setup, we denote $D_{s}$ and $D_{t}$ as source and target task datasets, respectively. Note that the formalization is equivalent to 1-shot prompting where the only provided example input-output pair comes from a different dataset. Table 14 of Appendix F presents illustrative examples of prompts in this setup. Our experiments suggest that $>$1-shot setup in cross-task prompting does not improve (and often deteriorates) performance.
@@
Examples of prompts are provided in Table 15 of Appendix F.
@@ 3 Results and Analysis
Datasets and experimental setup. Our corpus of tasks consists of ten source and five target tasks. We consider ARC-Easy Clark et al. (2018), AG-news Zhang et al. (2015), BoolQ Clark et al. (2019), Commonsense-QA Talmor et al. (2019), Conll2003-POS Tjong Kim Sang and De Meulder (2003), Conll2003-NER Tjong Kim Sang and De Meulder (2003), MNLI Williams et al. (2018), QQP Sharma et al. (2019), RACE Lai et al. (2017) and SST2 Socher et al. (2013) to be our source tasks. Our motivation remains to incorporate domain diversity and difficulty of the problems while choosing target tasks to emulate the “novel task” phenomenon as closely as possible since learning from cross-task examples makes sense only when the target task is truly data-scarce. Our target tasks are Social-i-QA Sap et al. (2019), SciQ Auer et al. (2023), MedMCQA Pal et al. (2022), Financial-Phrasebank Malo et al. (2014), and ARC-Challenge Clark et al. (2018); first four of these require domain expertise, while the last one is a more challenging version of the one present in the source tasks. In total, this gives us $50$ unique cross-task setups (see Appendix A for dataset details. Table 16 and Table 17 of Appendix F show task definitions corresponding to source and target tasks, respectively). The dataset size is standardised by sampling $10,000$ and $500$ examples for each source and target task, i.e., $|D_{s}|=10,000$ and $|D_{t}|=500$. For our experiments, we used the 7-billion and 13-billion variants of LLaMA-2 Touvron et al. (2023), and text-davinci-003 Brown et al. (2020) referred to as GPT3.5. We experiment with greedy and force decoding setup and selected greedy decoding as our standard for all experiments (see Appendix D for more details). We set the number of examples used to create cross-task context as one for all our experiments, unless mentioned explicitly.
@@
On average, cross-task prompting improves performance by 107% for LLaMA-2 7B, 18.6% for LLaMA-2 13B and 3.2% for GPT3.5. This is a strong argument for prompting in a cross-task manner when we lack labeled target data, especially using small models, which have poor zero-shot abilities  (Wei et al., 2022). We also experiment with LLaMA-2 7B Chat model to analyse the performance of cross-task prompting for instruction-tuned models (see Appendix C).
Performance with dissimilar source tasks. Few of the chosen source tasks have slight overlap in their broad topic or format with some of the target tasks, for example, source datasets like ARC-Easy, Commonsense-QA, RACE consists of questions in multiple-choice format, similar to target datasets like ARC-Challenge, Social-i-QA, SciQ and MedMCQA. To analyse the efficacy of cross-task prompting with completely dissimilar source tasks, let us consider only the source tasks whose topic and format are absolutely disjoint from all target tasks. The following five source tasks can be considered dissimilar to the target tasks – BoolQ, Conll2003-POS, Conll2003-NER, MNLI and QQP. As reported in Table 1, even with these dissimilar source tasks, we observe a substantial performance improvement of 80.5% for LLaMA-2 7B, 11.4% for LLaMA-2 13B, and 0.5% for GPT3.5 on average over zero-shot prompting. Moreover, if we consider only BoolQ, MNLI and QQP, the average performance gains become 105.9% for LLaMA-2 7B, 22.8% for LLaMA-2 13B, and 4% for GPT3.5. This shows that cross-task prompting can especially enable the relatively smaller LMs to substantially boost their zero-shot performance even using tasks dissimilar to the target.
Importance of source task definitions. To further investigate the role of task definitions in cross-task prompting, we check the effect of removing source task definitions on performance. We note an average drop of 11% for LLaMA-2 7B and 8% for LLaMA-2 13B when we prompt without using source task definitions (see Table 13 in Appendix). Hence, definitions play a crucial role in cross-task prompting.
@@
Semantically similar vs random example selection. As shown in Table 2, in a cross-task setup, choosing the examples randomly leads to substantially poorer performance than when we generate the context $C$ with semantically similar examples; concerning is the fact that, in many cases, it causes 0% accuracy. This may be caused by the model getting confused without any semantic alignment in the prompt. Furthermore, random labeling of in-context examples (Wei et al., 2023) result in near-random performance (see Table 12 in Appendix E).
@@ Appendix B Semantic context creation
Appendix C Cross-task prompting in instruction-tuned models
The performance gain with cross-task prompting for instruction-tuned models is lesser compared to that for the corresponding base model, due to the improvement in zero-shot performance of instruction-tuned models. For instance, Table 9 shows the performance for LLaMA-2 7B Chat model. We observe that it’s zero-shot performance is much better compared to the LLaMA-2 7B base model.
TAR SRC Zero-shot ARC-Easy AG-news BoolQ Com-QA C-POS C-NER MNLI QQP RACE SST2 LLaMA-2 7B Chat ARC-Challenge 42.00 51.00 39.20 45.00 51.40 37.00 30.20 37.00 37.20 51.40 39.00 Financial-Phrasebank 79.64 76.05 83.63 77.84 68.66 43.91 12.18 77.25 80.84 78.84 83.43 MedMCQA 30.60 31.00 30.20 31.60 34.60 30.00 26.80 30.20 27.40 31.00 30.20 SciQ 65.60 66.60 59.00 67.00 71.80 58.20 45.00 53.80 60.80 69.20 59.60 Social-i-QA 52.10 48.10 47.70 47.70 49.50 47.90 36.33 42.32 41.92 53.69 45.91
Table 9: Accuracy for cross-task setup, using one source example, with the LLaMA-2 7B Chat model. (Abbreviations: Com-QA $\rightarrow$ Commonsense-QA, C-POS $\rightarrow$ Conll2003-POS, C-NER $\rightarrow$ Conll2003-NER).
Appendix D Force decoding
@@
Appendix E Random labeling
@@
Appendix F Prompt details
@@
Appendix G Error analysis
@@
Appendix H Activation analysis on LLaMA-2 7B
Full correlation analysis is presented in Table 19.